# DocCGen: Document-based Controlled Code Generation

### v1

@@ Abstract
Figure 1: Illustration of shortcomings with fine-tuning and DocPrompting (zhou2022doccoder) approaches with an example for (a) NL to Bash task (uses GPT Neo 1.3B) and (b) NL to Ansible-YAML task (uses StarCoder2 3B) and the proposed DocCGen method to overcome the limitations.
@@ 1 Introduction
The Natural Language to Code (NL-to-Code) task has become pivotal in the intersection of natural language processing and programming. NL-to-Code systems can help engineers write a program efficiently by conveying their intentions at a higher level, as shown in Figure 1. Systems like Amazon code Whisperer [^2] https://aws.amazon.com/codewhisperer/, GitHub Co-pilot [^3] https://github.com/features/copilot/ perform well in NL-to-Code task due to large language models (LLM) trained on extensive data. While they perform well in general resource-rich languages like C++, Python, or Java, their practical usage in structured DSL is limited. DSLs are enterprise-specific languages with specialized schemas and syntax suitable for a specific domain or application [^4] https://w.wiki/6jCH. Numerous enterprises use structured languages like Bash, YAML, JSON and HCL (HashiCorp Configuration Language) with specific customizations for automation and to configure and manage infrastructure in IT environments. These languages or their customizations are potentially unseen by LMs during pre-training, limiting their practical usage (zan-etal-2022-language). Some existing methods attempt to address this challenge via in-context learning through examples (poesia2022synchromesh), by fine-tuning (10247987) or by using relevant documentation as additional context (zan-etal-2022-language; zhou2022doccoder; parvez2021retrieval; lu2022reacc). However, relevant context or samples available for DSL are often insufficient to incorporate diverse library schema rules or specialized structure knowledge in the LM (zan-etal-2022-language; wang2024grammar). This results in hallucination and different syntactic and semantic errors, as shown in Figure 1. However, enterprises usually maintain detailed documentation of their custom libraries (e.g. ansible modules, bash utilities), including the descriptions, schema, and syntax, to assist developers in enforcing structure and maintaining data integrity. We believe such schema and documentation can be better leveraged during code generation. Therefore, we propose a framework DocCGen that treats the NL-to-Code task as a two-step process, each heavily relying on the documentation. The first step identifies relevant code libraries for the task by retrieving the library documentation relevant to the NL query. The second step employs constrained decoding (CD) to guide code generation by using the grammar and schema rules extracted from the documentation of libraries identified in the first step, as shown in Figure 2. We evaluate this approach for two diverse and complex structured languages, Ansible YAML and Bash command. Generation for these languages is tricky due to complexities like the diverse library schemas, optional and required fields, the order-agnostic nature of fields, and inter-field dependencies. We believe studying these complex structures encompasses most of the challenges in other structured DSLs and allows easily extending DocCGen to other domains. Since the major challenge in DSLs is the limited availability of samples, we focus on enhancing performance for unseen code libraries or libraries with very few samples in the training corpus. Hence, we evaluate our approach in two settings: In-domain and Out-of-domain. Similar to zhou2022doccoder, none of the libraries in the test set are seen during training in the OOD setting. In the ID setting, every library in the test set has very few NL-to-Code pairs in the train set. DocCGen consistently improve over state-of-the-art models and techniques by a significant margin (Table 1, 2) across multiple settings.
@@
Finally, we introduce first publicly available benchmark dataset for NL to structured code generation task consisting of Ansible-YAML language. Intricate challenges in Ansible-YAML generation, like the complex structure and diverse module schemas, lead to subpar performance even for fine-tuned code LMs (Table 1). We curate NL to Ansible-YAML dataset with $18k$ samples with code snippets from more than $2500$ modules under OOD and ID settings (Table 5). More information and examples for Ansible-YAML are presented in section A.1. Besides this, we augment new NL to Ansible-YAML and existing NL to Bash dataset TLDR (zhou2022doccoder) with descriptions, detailed schema and grammar information from each library. We believe these datasets will advance research in constrained generation and handling low-resource or unseen data scenarios in structured DSLs.
@@ 2 Related Work
Constrained decoding: Controlled code generation using constraints has been previously studied majorly for the text-to-SQL task, using plan-based static templates (bhaskar2023benchmarking) or SQL parser-based semantic checks (scholak-etal-2021-picard). The database schema is fixed and given as input with a text query for text-to-SQL. However, we target a more complex problem involving multiple libraries and diverse schemas and use library documentation to solve this. poesia2022synchromesh and wang2024grammar use in-context learning via relevant samples or grammar strings and constrain the decoding further. However, in-context learning does not solve the issue of the correctness of the library. Hence, we instead follow a two-step process using library documentation. agrawal2023guiding uses constrained decoding for general-purpose languages like Java and C# using suggestions from intelligent parsers. However, such advanced parsers are uncommon for DSLs and might provide incomplete constraints. Hence, we use rules extracted from documentation more commonly available.
Context Based Controlled Generation like RAGs: Many existing methods retrieve the relevant context and augment it with the input prompt to improve the code generation (lu2022reacc; zan-etal-2022-language; zhou2022doccoder; parvez2021retrieval; ding2022cocomic). Although effective, these methods do not ensure schema and grammar adherence, especially for unseen libraries and languages. zhang2023repocoder and zan-etal-2022-language improve over vanilla retrieval-augmented code generation but require either architectural changes or extra pre-training. Hence, unlike these methods, we guide the generation by adjusting the output logits.
@@ 3.3.1 Sparse retrieval
We use the BM25 retrieval system Robertson1976RelevanceWO that uses sparse features such as word frequencies to calculate similarity with documents.
@@ 3.3.2 Dense retrieval
For dense retrieval systems, we fine-tune pre-trained ColBERTv2 santhanam2021colbertv2 and also use it in the zero-shot setting. Finally, we use the best results for the downstream generation task.
Training: We fine-tune ColBERTv2 based on triplet formed as $<q,D^{+},D^{-}>$. $D^{+}$ is the document of the libraries relevant to query $q$. $D^{-}$ is a set of documents of libraries that are not relevant to $q$ but are similar to $D^{+}$. For $q$ we prepare the training set as ($q$, $d_{1}^{+}$, $d_{2}^{+}$,…..,$d_{m}^{+}$, $d_{1}^{-}$, $d_{2}^{-}$….., $d_{n}^{-}$) where $d_{i}^{+}$ is the positive document, and each $d_{i}^{-}$ is a negative document which is not relevant to $q$. We select $n$ hard negatives using miniLM sentence BERT similarity scores similar to santhanam2021colbertv2. Using such a train set, we train ColBERTv2 by minimizing the distance between $q$ and $D^{+}$ and maximizing the distance between $q$ and $D^{-}$.
@@ 3.4 Constrained generation
String selection algorithm is used to constrain the model to generate exactly one string from a set of strings ($S$) {$s_{1},s_{2},s_{3}...,s_{n}$} (agrawal2023guiding). Initially, all the strings are tokenized, and we limit the vocabulary $V$ of the model to a set of tokens $t\in V$, which form the prefix of any string in $S$. Once a token $t_{i}$ among $t$ is sampled, all the strings that do not have $t_{i}$ as a prefix are discarded. The same process is repeated until exactly one string is chosen.
@@ 4.2 Bash command
Since we primarily focus on improving performance for unseen libraries and low-resource data settings, we select the TLDR (zhou2022doccoder) as our primary dataset for NL to Bash. TLDR consists of $1503$ bash utilities across the train and test samples. This data consists of $7342$ NL to bash pairs with $4.3$ pairs for every utility. Train and test splits of this data consist of $7342$ NL to bash pairs. A low number of samples for each utility creates a scarce data scenario.
Other than this, we also use NL2Bash (lin2018nl2bash) dataset consisting of $8090$ train and $609$ test samples for $100$ bash utilities. Due to the high number of NL to bash pairs for every bash utility, this dataset allows us to check performance for resource-rich settings. However, Since this is not the major focus of the work, results for NL2Bash are included in Appendix (Table 11)
@@ 5.2 Baselines
Existing methods like APICoder (zan-etal-2022-language) pre-train models on abundant documentation and code samples for general-purpose languages like Python. Replicating this setup for structured DSLs is challenging due to the scarcity of available code samples. Hence, for best comparison, we pre-train our models on Linux man pages for bash and Ansible documentation for YAML, ensuring no data leakage from fine-tuning datasets. We then fine-tune the pre-trained model on respective NL-to-Code tasks and compare its performance with DocCGen. We also perform ablation studies with Base + IR setup for the pre-trained models (Table 9, 10). Details of pre-training data are provided in the Appendix (section A.4, A.2).
@@
We adopt DocPrompting (zhou2022doccoder) as a baseline for OOD split through the TLDR dataset because it is a RAG-based approach, currently state-of-the-art for TLDR. Additionally, Unlike other RAG-based methods (parvez2021retrieval; zhang2023repocoder), it uses documentation instead of abundant code samples, aligning better with our DSL use case with scarce examples.
@@ 5.3 Models
Generator We include different sized state-of-the-art code language models in our evaluation, including StarCoder2 family (3B, 7B, 15B) (lozhkov2024starcoder), and CodeLlama 34B (roziere2023code). Due to resource constraints to fine-tune large parameter models like CodeLlama 34B and Starcoder2 15B, we experiment with their instruction-tuned version in a 3-shot setting and present their results in Appendix (Table 6). Further, our evaluation includes a fine-tuned GPT Neo 1.3B (gpt-neo) version to compare with the DocPrompting baseline. We use beam search inference decoding for all the base fine-tuned models with beam width $5$.
@@ 5.4 Evaluation metrics
Bash command: Evaluation metrics for bash include 1) Command name accuracy (CMD Acc): This metric evaluates the exact match of bash utility in the command (e.g. tar, cat). 2) Exact Match: Exact match of full generated command and reference command 3) Token F1 score (zhou2022doccoder).
Ansible YAML: We leverage $2$ evaluation metrics from 10247987 - Schema Correct, and Ansible Aware. Additionally, we introduce the Module Acc metric, which measures the correctness of the generated YAML module. This metric is similar to the CMD Acc metric in bash. Refer to A.1.6 for a detailed description of metrics.
@@ 6 Results and Analysis
We observe that extended pre-training does not improve performance in structured DSLs with limited code samples in the documentation. Therefore, we use an IR-based approach that focuses on retrieving utility descriptions, unlike zhou2022doccoder, which retrieves passages with options (flags and sub-commands) and utilities. This targeted detection reduces the search space for IR from $400k$ to $1.5k$ documents, leading to a notable improvement in Hits@1 (Table 4). This improves CMD Acc from $27.59\%$ to $38.32\%$ when the model is constrained to follow the Hits@1 retrieved library template (Table 3). CMD Acc consistently improves for the ID setting by around $6\%$ to $12\%$ (Table 3). For YAML, Module Acc significantly improves compared to the fine-tuned baselines, especially in the OOD setting ($\sim 10\%$). Further, we restrict the model to follow one of the templates for $k$ retrieved libraries. CMD Acc and Module Acc drop with a higher value of $k$ (Table 7, 8), which is expected since relaxing constraints on the model tend to approach its performance towards the baselines.
@@ 7 Conclusion
We propose DocCGen, a novel framework for NL-to-Code generation for structured DSLs. DocCGen decomposes the NL-to-Code generation into two steps involving the detection of relevant libraries in the first step and using schema and grammar rules extracted from the documentation of these libraries to guide the decoding in the second step. We evaluate the performance of DocCGen for two complex structured languages, Bash command and Ansible YAML, involving two settings, OOD and ID. Our approach outperforms state-of-the-art techniques consistently across all metrics for different-sized models. It reduces syntactic and semantic errors in code, particularly for unseen libraries and low-resource data settings. We also contribute the first publicly available benchmark dataset for NL to Ansible-YAML task. We augment NL to Ansible-YAML and TLDR dataset with description and schema information. We hope this work will help advance research in solving DSL related tasks and constrained generation.
@@
We break down code generation in to two steps: a) Information Retrieval and b) Generation based on retrieved documentation. Therefore, errors in retrieval for the user query may cascade to the generation step. Even though, we see that leveraging documentation in this pipeline-based approach results in significant improvements for custom settings, we believe that jointly training the retriever and generator might mitigate these errors. This can be explored as a part of future work. Apart from this, constrained decoding adds a computational overhead during inference. However, since we add the rules on top of efficient greedy decoding, constrained decoding is practical to use as beam search decoding which is widely adopted is similarly computationally heavy. Still, this can be mitigated using constrained generation in speculative decoding similar to wang2024grammar. Such improvements can easily be integrated with our framework. Further, parser-based methods to automatically integrate grammar rules during decoding can help generalize DocCGen to a larger scale.
@@ A.5 Hyperparameter details
We use NVIDIA A100 $80$ GB GPUs to perform inference and training for all the experiments. We use the standard HuggingFace transformers wolf-etal-2020-transformers with accelerate to load, train, and perform inference for all the models. For constrained decoding we use HuggingFace logits processor [^9] https://huggingface.co/docs/transformers/en/internal/generation_utils#transformers.LogitsProcessor.

### v2

@@ Abstract
Figure 1: Illustration of shortcomings with fine-tuning and DocPrompting (Zhou et al., 2022) approaches with an example for (a) NL to Bash task (uses GPT Neo 1.3B) and (b) NL to Ansible-YAML task (uses StarCoder2 3B) and the proposed DocCGen method to overcome the limitations.
@@ 1 Introduction
The Natural Language to Code (NL-to-Code) task has become pivotal in the intersection of natural language processing and programming. NL-to-Code systems can help engineers write a program efficiently by conveying their intentions at a higher level, as shown in Figure 1. Systems like Amazon code Whisperer [^2] https://aws.amazon.com/codewhisperer/, GitHub Co-pilot [^3] https://github.com/features/copilot/ perform well in NL-to-Code task due to large language models (LLM) trained on extensive data. While they perform well in general resource-rich languages like C++, Python, or Java, their practical usage in structured DSL is limited. DSLs are enterprise-specific languages with specialized schemas and syntax suitable for a specific domain or application [^4] https://w.wiki/6jCH. Numerous enterprises use structured languages like Bash, YAML, JSON and HCL (HashiCorp Configuration Language) with specific customizations for automation and to configure and manage infrastructure in IT environments. These languages or their customizations are potentially unseen by LMs during pre-training, limiting their practical usage (Zan et al., 2022). Some existing methods attempt to address this challenge via in-context learning through examples (Poesia et al., 2022), by fine-tuning (Pujar et al., 2023) or by using relevant documentation as additional context (Zan et al., 2022; Zhou et al., 2022; Parvez et al., 2021; Lu et al., 2022). However, relevant context or samples available for DSL are often insufficient to incorporate diverse library schema rules or specialized structure knowledge in the LM (Zan et al., 2022; Wang et al., 2024). This results in hallucination and different syntactic and semantic errors, as shown in Figure 1. However, enterprises usually maintain detailed documentation of their custom libraries (e.g. ansible modules, bash utilities), including the descriptions, schema, and syntax, to assist developers in enforcing structure and maintaining data integrity. We believe such schema and documentation can be better leveraged during code generation. Therefore, we propose a framework DocCGen that treats the NL-to-Code task as a two-step process, each heavily relying on the documentation. The first step identifies relevant code libraries for the task by retrieving the library documentation relevant to the NL query. The second step employs constrained decoding (CD) to guide code generation by using the grammar and schema rules extracted from the documentation of libraries identified in the first step, as shown in Figure 2. We evaluate this approach for two diverse and complex structured languages, Ansible YAML and Bash command. Generation for these languages is tricky due to complexities like the diverse library schemas, optional and required fields, the order-agnostic nature of fields, and inter-field dependencies. We believe studying these complex structures encompasses most of the challenges in other structured DSLs and allows easily extending DocCGen to other domains. Since the major challenge in DSLs is the limited availability of samples, we focus on enhancing performance for unseen code libraries or libraries with very few samples in the training corpus. Hence, we evaluate our approach in two settings: In-domain and Out-of-domain. Similar to Zhou et al. (2022), none of the libraries in the test set are seen during training in the OOD setting. In the ID setting, every library in the test set has very few NL-to-Code pairs in the train set. DocCGen consistently improve over state-of-the-art models and techniques by a significant margin (Table 1, 2) across multiple settings.
@@
Finally, we introduce first publicly available benchmark dataset for NL to structured code generation task consisting of Ansible-YAML language. Intricate challenges in Ansible-YAML generation, like the complex structure and diverse module schemas, lead to subpar performance even for fine-tuned code LMs (Table 1). We curate NL to Ansible-YAML dataset with $18k$ samples with code snippets from more than $2500$ modules under OOD and ID settings (Table 5). More information and examples for Ansible-YAML are presented in section A.1. Besides this, we augment new NL to Ansible-YAML and existing NL to Bash dataset TLDR (Zhou et al., 2022) with descriptions, detailed schema and grammar information from each library. We believe these datasets will advance research in constrained generation and handling low-resource or unseen data scenarios in structured DSLs.
@@ 2 Related Work
Constrained decoding: Controlled code generation using constraints has been previously studied majorly for the text-to-SQL task, using plan-based static templates (Bhaskar et al., 2023) or SQL parser-based semantic checks (Scholak et al., 2021). The database schema is fixed and given as input with a text query for text-to-SQL. However, we target a more complex problem involving multiple libraries and diverse schemas and use library documentation to solve this. Poesia et al. (2022) and Wang et al. (2024) use in-context learning via relevant samples or grammar strings and constrain the decoding further. However, in-context learning does not solve the issue of the correctness of the library. Hence, we instead follow a two-step process using library documentation. Agrawal et al. (2023) uses constrained decoding for general-purpose languages like Java and C# using suggestions from intelligent parsers. However, such advanced parsers are uncommon for DSLs and might provide incomplete constraints. Hence, we use rules extracted from documentation more commonly available.
Context Based Controlled Generation like RAGs: Many existing methods retrieve the relevant context and augment it with the input prompt to improve the code generation (Lu et al., 2022; Zan et al., 2022; Zhou et al., 2022; Parvez et al., 2021; Ding et al., 2022). Although effective, these methods do not ensure schema and grammar adherence, especially for unseen libraries and languages. Zhang et al. (2023) and Zan et al. (2022) improve over vanilla retrieval-augmented code generation but require either architectural changes or extra pre-training. Hence, unlike these methods, we guide the generation by adjusting the output logits.
@@ 3.3.1 Sparse retrieval
We use the BM25 retrieval system Robertson and Jones (1976) that uses sparse features such as word frequencies to calculate similarity with documents.
@@ 3.3.2 Dense retrieval
For dense retrieval systems, we fine-tune pre-trained ColBERTv2 Santhanam et al. (2021) and also use it in the zero-shot setting. Finally, we use the best results for the downstream generation task.
Training: We fine-tune ColBERTv2 based on triplet formed as $<q,D^{+},D^{-}>$. $D^{+}$ is the document of the libraries relevant to query $q$. $D^{-}$ is a set of documents of libraries that are not relevant to $q$ but are similar to $D^{+}$. For $q$ we prepare the training set as ($q$, $d_{1}^{+}$, $d_{2}^{+}$,…..,$d_{m}^{+}$, $d_{1}^{-}$, $d_{2}^{-}$….., $d_{n}^{-}$) where $d_{i}^{+}$ is the positive document, and each $d_{i}^{-}$ is a negative document which is not relevant to $q$. We select $n$ hard negatives using miniLM sentence BERT similarity scores similar to Santhanam et al. (2021). Using such a train set, we train ColBERTv2 by minimizing the distance between $q$ and $D^{+}$ and maximizing the distance between $q$ and $D^{-}$.
@@ 3.4 Constrained generation
String selection algorithm is used to constrain the model to generate exactly one string from a set of strings ($S$) {$s_{1},s_{2},s_{3}...,s_{n}$} (Agrawal et al., 2023). Initially, all the strings are tokenized, and we limit the vocabulary $V$ of the model to a set of tokens $t\in V$, which form the prefix of any string in $S$. Once a token $t_{i}$ among $t$ is sampled, all the strings that do not have $t_{i}$ as a prefix are discarded. The same process is repeated until exactly one string is chosen.
@@ 4.2 Bash command
Since we primarily focus on improving performance for unseen libraries and low-resource data settings, we select the TLDR (Zhou et al., 2022) as our primary dataset for NL to Bash. TLDR consists of $1503$ bash utilities across the train and test samples. This data consists of $7342$ NL to bash pairs with $4.3$ pairs for every utility. Train and test splits of this data consist of $7342$ NL to bash pairs. A low number of samples for each utility creates a scarce data scenario.
Other than this, we also use NL2Bash (Lin et al., 2018) dataset consisting of $8090$ train and $609$ test samples for $100$ bash utilities. Due to the high number of NL to bash pairs for every bash utility, this dataset allows us to check performance for resource-rich settings. However, Since this is not the major focus of the work, results for NL2Bash are included in Appendix (Table 11)
@@ 5.2 Baselines
Existing methods like APICoder (Zan et al., 2022) pre-train models on abundant documentation and code samples for general-purpose languages like Python. Replicating this setup for structured DSLs is challenging due to the scarcity of available code samples. Hence, for best comparison, we pre-train our models on Linux man pages for bash and Ansible documentation for YAML, ensuring no data leakage from fine-tuning datasets. We then fine-tune the pre-trained model on respective NL-to-Code tasks and compare its performance with DocCGen. We also perform ablation studies with Base + IR setup for the pre-trained models (Table 9, 10). Details of pre-training data are provided in the Appendix (section A.4, A.2).
@@
We adopt DocPrompting (Zhou et al., 2022) as a baseline for OOD split through the TLDR dataset because it is a RAG-based approach, currently state-of-the-art for TLDR. Additionally, Unlike other RAG-based methods (Parvez et al., 2021; Zhang et al., 2023), it uses documentation instead of abundant code samples, aligning better with our DSL use case with scarce examples.
@@ 5.3 Models
Generator We include different sized state-of-the-art code language models in our evaluation, including StarCoder2 family (3B, 7B, 15B) (Lozhkov et al., 2024), and CodeLlama 34B (Roziere et al., 2023). Due to resource constraints to fine-tune large parameter models like CodeLlama 34B and Starcoder2 15B, we experiment with their instruction-tuned version in a 3-shot setting and present their results in Appendix (Table 6). Further, our evaluation includes a fine-tuned GPT Neo 1.3B (Black et al., 2021) version to compare with the DocPrompting baseline. We use beam search inference decoding for all the base fine-tuned models with beam width $5$.
@@ 5.4 Evaluation metrics
Bash command: Evaluation metrics for bash include 1) Command name accuracy (CMD Acc): This metric evaluates the exact match of bash utility in the command (e.g. tar, cat). 2) Exact Match: Exact match of full generated command and reference command 3) Token F1 score (Zhou et al., 2022).
Ansible YAML: We leverage $2$ evaluation metrics from Pujar et al. (2023) - Schema Correct, and Ansible Aware. Additionally, we introduce the Module Acc metric, which measures the correctness of the generated YAML module. This metric is similar to the CMD Acc metric in bash. Refer to A.1.6 for a detailed description of metrics.
@@ 6 Results and Analysis
We observe that extended pre-training does not improve performance in structured DSLs with limited code samples in the documentation. Therefore, we use an IR-based approach that focuses on retrieving utility descriptions, unlike Zhou et al. (2022), which retrieves passages with options (flags and sub-commands) and utilities. This targeted detection reduces the search space for IR from $400k$ to $1.5k$ documents, leading to a notable improvement in Hits@1 (Table 4). This improves CMD Acc from $27.59\%$ to $38.32\%$ when the model is constrained to follow the Hits@1 retrieved library template (Table 3). CMD Acc consistently improves for the ID setting by around $6\%$ to $12\%$ (Table 3). For YAML, Module Acc significantly improves compared to the fine-tuned baselines, especially in the OOD setting ($\sim 10\%$). Further, we restrict the model to follow one of the templates for $k$ retrieved libraries. CMD Acc and Module Acc drop with a higher value of $k$ (Table 7, 8), which is expected since relaxing constraints on the model tend to approach its performance towards the baselines.
@@ 7 Conclusion
We propose DocCGen, a novel framework for NL-to-Code generation for structured DSLs. DocCGen decomposes the NL-to-Code generation into two steps involving the detection of relevant libraries in the first step and using schema and grammar rules extracted from the documentation of these libraries to guide the decoding in the second step. We evaluate the performance of DocCGen for two complex structured languages, Bash command and Ansible YAML, involving two settings, OOD and ID. Our approach outperforms state-of-the-art techniques consistently across all metrics for different-sized models. It reduces syntactic and semantic errors in code, particularly for unseen libraries and low-resource data settings. We also contribute the first publicly available benchmark dataset for NL to Ansible-YAML task. We augment NL to Ansible-YAML and TLDR dataset with description and schema information. We hope this work will help advance research in solving DSL-related tasks and constrained generation.
@@
We break down code generation in to two steps: a) Information Retrieval and b) Generation based on retrieved documentation. Therefore, errors in retrieval for the user query may cascade to the generation step. Even though, we see that leveraging documentation in this pipeline-based approach results in significant improvements for custom settings, we believe that jointly training the retriever and generator might mitigate these errors. This can be explored as a part of future work. Apart from this, constrained decoding adds a computational overhead during inference. However, since we add the rules on top of efficient greedy decoding, constrained decoding is practical to use as beam search decoding which is widely adopted is similarly computationally heavy. Still, this can be mitigated using constrained generation in speculative decoding similar to Wang et al. (2024). Such improvements can easily be integrated with our framework. Further, parser-based methods to automatically integrate grammar rules during decoding can help generalize DocCGen to a larger scale.
@@ A.5 Hyperparameter details
We use NVIDIA A100 $80$ GB GPUs to perform inference and training for all the experiments. We use the standard HuggingFace transformers Wolf et al. (2020) with accelerate to load, train, and perform inference for all the models. For constrained decoding we use HuggingFace logits processor [^9] https://huggingface.co/docs/transformers/en/internal/generation_utils#transformers.LogitsProcessor.